how many seconds to cover 17 miles at 53 miles per hour
Answer by Alan3354(69443)

Here's an example. Do it like this:
driving at 55 mph one covers 19 miles in how many seconds
----------
d = rt
t = d/r
t = 19/55 hours
t = 3600*19/55 seconds
-----------------------
If you have a lot of these, make an Excel sheet.